OpenStudy (kevon18):

Find the function from the given zeros. 2, 3, and i

OpenStudy (sshayer):

i tried to post but every time it is posted in other so i write in detail. if a,b,c+id are the zeros. then zeros are a,b,c+id,c-id \[f(x)=(x-a)(x-b)\left\{ x-\left( c+id \right) \right\} ~\left\{ x-\left( c-id \right) \right\}\] \[=\left( x-a \right)\left( x-b \right)\left\{ \left( x-c \right)-\iota d \right\}\left\{ \left( x-c \right)+\iota d \right\}\] \[=\left( x-a \right)\left( x-b \right)\left\{ \left( x-c \right)^2-\iota ^2d^2 \right\}\] \[put~\iota^2=-1\]

OpenStudy (catpidigm):

wow that's a lot

OpenStudy (bonnieisflash1.0):

need help

OpenStudy (dumbcow):

\[x = 2 \rightarrow x-2 =0\] \[x=3 \rightarrow x-3 =0\] \[x=i \rightarrow x^2 = i^2 = -1 \rightarrow x^2 +1 = 0\] \[f(x) = (x-2)(x-3)(x^2+1)\]